OpenStudy (amandabonfiglio):

Derive the equation of the parabola with a focus at (−5, 5) and a directrix of y = −1.

OpenStudy (idkanything):

@YanaSidlinskiy

OpenStudy (idkanything):

WELCOME TO OPENSTUDY!!!!

OpenStudy (amandabonfiglio):

i have the answer choices! and thank you :)

OpenStudy (idkanything):

ok what are they?

OpenStudy (amandabonfiglio):

f(x)=-1/12(x-5)^2+2 f(x)=1/12(x-5)^2+2 f(x)=-1/12(x+5)^2+2 f(x)=1/12(x+5)^2+2

OpenStudy (idkanything):

yana pls help

OpenStudy (yanasidlinskiy):

It'd D.